There are n arithmetic means between 4 and 31 such that the second mean :last mean=5:14.find n.

Grade:11

1 Answers

Bhupesh Kashinath Kocharekar
5 years ago
Nequals 8
D equals 3 a=4 total no of terms in ap n+2
4th = a + (n+2-1)d=34
*n+1)d=31-4=27
(31-d)÷(4+d)=14÷5 solve